# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

### v3

@@ 3.2 No-Thinking-Med-R1
  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound.
  • Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography
  • Micro - Microscopy Images; X-Ray - X-Ray Imaging
@@ 4 Experiment & Results
This section introduces our experimental setup, implementation details, and results.
@@ 4.1 Setup
Implementation Details Training is conducted on HGX H100 [7] server with 2×H100 GPUs (80GB VRAM) using PyTorch [27] and FlashAttention-2 [11] for optimized efficiency. We initialize from Qwen2-VL-2B-Instruct [2] with full parameter tuning, employing per-GPU batch size 1 (effective batch size 4 via 2-step gradient accumulation) and bfloat16 mixed precision. Input sequences combine visual embeddings from 328×328 resolution images (max 401k pixels) with textual prompts truncated to 1,024 tokens. The GRPO policy generates four candidate rationales per sample, with a sampling temperature of $\tau=0.7$. Each training configuration is run for one epoch.
@@
  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the remaining 7 unseen modalities.
  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the remaining four.
@@
Table 2: Performance Comparison of VLMs on Eight Modality: Performance comparison across eight medical imaging modalities demonstrates that our GRPO-finetuned model surpasses zero-shot general-purpose VLMs, domain-specific medical VLMs, and conventional supervised fine-tuning approaches, achieving superior accuracy while maintaining scalability for clinical deployment. Meanwhile, in each column, the best and second-best performances are marked in red and blue. Naming details are as follows: CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging.
Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
Zero-shot VLM
[…]
Qwen2-VL-2B [2] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
Qwen2-VL-7B [2] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
Zero-shot Medical VLM
[…]
Fine-tuned VLM
Med-R1-2B (GRPO) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
@@ 4.2 Cross-Modality Generalization
We comprehensively evaluate Med-R1’s adaptability across 8 medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging. Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
Results on generalization. To evaluate Med-R1’s cross-modality generalization, we measure its accuracy across eight distinct medical imaging modalities (Table 1). The Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability. Med-R1 achieves a strong overall accuracy of 69.91%, demonstrating its ability to generalize across diverse medical imaging modalities. Notably, models trained on CT, MRI, and X-Ray exhibit the highest generalization capability, with overall scores of 71.44%, 71.26%, and 72.35%, respectively. In contrast, models trained on Fundus Photography and Microscopy images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains. Importantly, the overall test accuracy of 69.91% highlights Med-R1’s ability to perform well across unseen imaging modalities, despite being trained on a single domain at a time. This result underscores the effectiveness of reinforcement learning in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five medical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
@@
4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across 5 important medical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five medical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs. GRPO). Results demonstrate that our compact GRPO-tuned 2B model surpasses even 36× larger zero-shot models and achieves stronger cross-task generalization compared to SFT. Meanwhile, in each column, the best and second-best performances are marked in red and blue.
@@ 4.3 Cross-Task Generalization
Zero-shot Medical VLM
@@
Med-R1-2B (GRPO) 62.88 66.08 65.87 98.24 80.14 74.64
@@ 5 Conclusion
In this work, we introduce Med-R1, a RL-enhanced VLM designed to improve medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization. Comprehensive evaluations across eight medical imaging modalities and five medical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
Despite its compact 2B parameter size, Med-R1 achieves competitive performance against 36×36$\times$36 × larger models, demonstrating that scalability does not have to come at the cost of efficiency. Its lightweight nature significantly lowers computational and deployment barriers, making it more practical for real-world clinical applications where resource constraints are critical.
While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions. Future research could further enhance Med-R1 by integrating additional multi-modal medical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.
6 Acknowledgement
This work is supported in part by the National Institutes of Health under award numbers R01CA272991, R01EB032680, R01DE033512, and U54CA274513.

### v4

@@ 3.1 Group Relative Policy Optimization (GRPO)
  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography.
  • OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging
Table 2: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.
No-Think RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 99.54 82.18 80.00 46.24 72.13 72.50 76.06 70.36 74.88
MRI 79.27 99.72 74.74 42.86 79.94 73.50 84.67 69.10 75.48
X-Ray 78.37 69.94 97.09 46.43 70.75 80.97 69.81 70.63 73.00
US 59.82 66.53 73.37 100.00 70.60 68.03 68.63 79.28 73.28
Der 56.34 67.94 73.87 46.14 98.00 69.67 67.45 67.57 68.37
FP 59.55 68.62 73.93 51.59 72.43 93.99 72.64 70.36 70.39
OCT 80.47 73.50 77.15 49.08 65.08 70.95 100.00 67.21 72.93
Micro 64.15 66.56 76.84 53.13 68.91 70.95 73.47 98.56 71.50
Overall 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.48
Reward design: We follow [12] and use two types of reward: format and accuracy. Firstly, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag. The format reward is designed to check if the aforementioned tags are present in the final response. A reward score of 1 will be given if they exist and are correct. This helps the model to organize its thoughts and answer in a structured format for the ease of reading. The accuracy reward is a rule-based reward that checks if the actual answer matches with the ground truth. Similarly, a reward score of 1 is given when the results match. In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
3.2 No-Thinking-Med-R1
Previous work [14] found that, in image classification tasks, removing both reasoning and format supervision during RL could sometimes improve performance. However, that finding is specific to structured classification tasks with simple output space, where reasoning is rarely required. In contrast, medical VQA involves multimodal, semantically complex inputs, and reasoning failures can arise from domain mismatch rather than verbosity alone [12, 6]. Given the difficulty of obtaining reliable CoT annotations in medical settings, we further investigate the role of intermediate reasoning in RL post-training. We revise the instruction prompt as {Question}. Output the single-letter choice (A, B, C, D,...) in <answer>...</answer> tags., where {Question} will be replaced by each specific question. By doing so, we encourage the model only to output the final answer without any explicit thinking process. The accuracy reward is maintained, where the reward score is 1 when the extracted answer matches the ground truth labels. Note that when there is any text outside the <answer> tag, i.e., an explicit thinking process exists, the extracted content will be null, and therefore the accuracy reward will be 0. Therefore, the model will be forced to generate only the answers.
Table 3: Performance comparison across eight medical modalities. Our GRPO-finetuned model outperforms zero-shot general-purpose VLMs, medical-domain VLMs, and supervised fine-tuning baselines, while maintaining scalability. The best and second-best performances per column are highlighted in red and blue. Modality abbreviations: CT – Computed Tomography; MRI – Magnetic Resonance Imaging; US – Ultrasound; Der – Dermoscopy; FP – Fundus Photography; OCT – Optical Coherence Tomography; Micro – Microscopy; X-Ray – X-ray Imaging.
Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
Zero-shot VLMs
[…]
$\text{LLaVA}^{\dagger}$ [28] 17.73 26.72 30.70 18.66 49.74 47.11 33.73 28.87 31.66
[…]
Qwen2-VL-2B [3] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
Qwen2-VL-7B [3] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
Qwen2.5-VL-3B [31] 53.87 54.23 61.84 32.69 52.94 62.47 56.23 59.64 54.24
Qwen2.5-VL-7B [31] 60.44 58.44 73.99 30.66 62.48 67.30 61.20 67.84 60.29
Qwen2.5-VL-72B [31] 66.18 68.74 77.59 49.81 69.75 71.04 69.22 69.37 67.71
Zero-shot Medical VLMs
[…]
$\text{RadFM}^{\dagger}$ [32] 27.56 24.06 30.95 16.57 39.21 36.89 32.80 27.97 29.50
[…]
Fine-tuned VLMs
Qwen2.5-VL-3B (SFT) 56.06 60.81 69.23 41.77 60.11 69.19 63.95 65.66 60.85
Qwen2-VL-2B (GRPO + Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
Qwen2-VL-2B (GRPO + Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
Qwen2.5-VL-3B (GRPO + Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
@@ 4.1 Setup
Implementation Details Training is conducted on HGX H100 [34] server with 2×H100 GPUs (80GB VRAM) using PyTorch [35] and FlashAttention-2 [36] for optimized efficiency. We initialize from Qwen2-VL-2B-Instruct [3] with full parameter tuning, employing per-GPU batch size 1 (effective batch size 4 via 2-step gradient accumulation) and bfloat16 mixed precision. Input sequences combine visual embeddings from 328×328 resolution images (max 401k pixels) with textual prompts truncated to 1,024 tokens. The GRPO policy generates four candidate rationales per sample, with a sampling temperature of $\tau=0.7$. Each training is run for one epoch.
@@
  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the other seven modalities.
  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the other four.
@@ 4.2 Cross-Modality Generalization
We comprehensively evaluate Med-R1’s adaptability across eight medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging. Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
Table 4: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition. The base model is Qwen2-VL-2B.
@@
Table 5: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. The base model is Qwen2-VL-2B.
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification 96.33 59.29 64.22 99.04 76.56 79.09
Disease Diagnosis 55.00 98.30 63.76 98.08 85.23 80.07
Lesion Grading 54.22 58.04 96.33 97.37 72.30 75.65
Modality Recognition 56.84 57.95 55.05 99.79 73.15 68.56
Other Attributes 56.29 58.00 52.29 97.90 99.29 72.75
Overall 63.74 66.32 66.33 98.44 81.31 75.22
Results on generalization. To evaluate Med-R1’s cross-modality generalization, we measure its accuracy across eight distinct medical imaging modalities (Table 1). Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability. Med-R1 achieves a strong overall accuracy of 69.91%, demonstrating its ability to generalize across diverse medical imaging modalities. Notably, models trained on CT, MRI, and X-Ray exhibit the highest generalization capability, with overall scores of 71.44%, 71.26%, and 72.35%, respectively. In contrast, models trained on Fundus Photography and Microscopy images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains. Importantly, the overall test accuracy of 69.91% highlights Med-R1’s ability to perform well across unseen imaging modalities, despite being trained on a single domain at a time. This result underscores the effectiveness of reinforcement learning in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
Table 6: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs. The best and second-best performances are marked in red and blue.
@@
Qwen2.5-VL-3B [31] 35.23 50.45 52.79 85.23 54.77 55.69
Qwen2.5-VL-7B [31] 41.00 61.32 54.13 97.78 70.45 64.94
Qwen2.5-VL-72B [31] 57.22 62.55 60.32 98.20 77.41 71.14
Zero-shot Medical VLMs
[…]
$\text{RadFM}^{\dagger}$ [32] 13.31 21.69 30.35 26.64 43.85 26.99
@@
Qwen2.5-VL-3B (SFT) 54.91 57.75 64.04 84.95 71.56 66.64
Qwen2-VL-2B (GRPO + Think) 62.88 66.08 65.87 98.24 80.14 74.64
Qwen2-VL-2B (GRPO + Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
Qwen2.5-VL-3B (GRPO + Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across five important clinical tasks [33]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
Figure 3: Cross-modality accuracy difference between No-Thinking-Med-R1 and Med-R1. Each cell shows the performance gap (%) when trained on one modality (y-axis) and evaluated on others (x-axis). Red indicates improvement; blue indicates degradation.
Figure 4: Cross-task accuracy difference between No-Thinking-Med-R1 and Med-R1. Each cell shows the performance gap (%) when trained on one task (y-axis) and tested on another (x-axis). Red indicates improvement; blue indicates degradation.
Figure 5: Rewards & Length. The top row shows the change in RL rewards during Med-R1 training across eight imaging modalities and five task types. We observe that the model quickly converges to stable reward values within 100–200 steps, indicating the computational efficiency of our RL post-training process. The bottom row shows the output completion length over training steps. Interestingly, the output length gradually decreases as training progresses, suggesting that the reward function discourages overly verbose or unfaithful reasoning. This observation challenges the conventional assumption that longer reasoning necessarily benefits vision-language performance in VQA.
4.4 Analysis of Thinking Process
We present the results of No-Thinking-Med-R1 in this section. Cross-task performance is shown in Table 5, and cross-modality performance is shown in Table 2. Overall, No-Thinking-Med-R1 achieves stronger in-domain performance across all settings compared to Med-R1, suggesting that removing explicit reasoning generation may lead to more effective task-specific learning. To assess generalization, we compare Med-R1 and No-Thinking-Med-R1 across modalities and tasks. As visualized in Figure 3 and Figure 4, No-Thinking-Med-R1 consistently outperforms Med-R1 in cross-modality settings. In cross-task scenarios, however, the results are more mixed—showing gains in some tasks and declines in others. These findings reinforce the practical motivation behind Med-R1. In clinical domains, acquiring high-quality CoT supervision is prohibitively costly and often impractical. While reasoning improves performance in general domains, our results show that this does not necessarily hold in medical settings, where unsupervised rationales may become unreliable under domain shift. Med-R1 demonstrates that even without reasoning supervision, RL can offer a robust and efficient path for adapting VLMs to medicine.
4.5 Analysis of Rewards and Completion Length.
The top row of Figure 5 shows the reward dynamics during GRPO-based reinforcement learning. Across all modalities and tasks, Med-R1 converges within 100–200 steps, as evidenced by the log-scale reward curves. This rapid and consistent convergence highlights the efficiency of our training pipeline and the model’s ability to quickly adapt to the reward structure. Moreover, the smooth reward trajectories suggest that GRPO remains stable even with limited RL updates, making it well-suited for low-resource scenarios.
Unlike in scientific domains such as mathematics [12], where model responses become increasingly detailed as thinking deepens, we observe a different trend in Med-RL training: the average response length decreases rapidly and then fluctuates (Figure 5). Rather than generating elaborate rationales, the model learns to produce shorter, more direct outputs—potentially favoring conciseness over verbosity. This aligns with our findings in subsection 4.4, suggesting that extended reasoning may not always benefit cross-domain VQA, especially under large domain shifts. We hypothesize that RL without CoT supervision encourages the model to rely on its pre-trained knowledge when generating rationales. While this may be effective in general domains, it can lead to hallucinated or misaligned reasoning in specialized domains like medical images. These results challenge the common assumption that more thinking always improves performance, and highlight that the “quality”—rather than the “presence”—of thinking is a crucial factor affecting the outcome of RL post-training.
4.6 Limitations and Future Work.
This work marks an initial step in applying RL to medical vision-language models. We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging. In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time. Future directions include extending Med-R1 to support multiple inputs. We also plan to incorporate patient context and explore more advanced reasoning frameworks for clinical deployment.
@@ 5 Conclusion
We present Med-R1, a RL-enhanced vision-language model that improves medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 surpasses the limitations of supervised fine-tuning, achieving strong cross-modality and cross-task generalization. Despite its compact 2B size, it matches or exceeds much larger models—including domain-specific and general-purpose VLMs—while remaining lightweight and deployable in resource-constrained clinical settings. Interestingly, we find that RL-based learning without explicit step-by-step rationales can outperform reasoning-heavy approaches, particularly under domain shift. This suggests that the quality, not quantity, of reasoning is key for robust medical adaptation. Med-R1 opens promising directions for developing scalable, reliable, and efficient AI systems in healthcare.